The height of babies are normally distributed with a mean of 65 centimeter and a standard deviation of 9 centimeter. find the percentage of the followong baby heights
- less than 54 centimeter
-at least 80 centimeter (more)
-between 70 and 86 centimer (between)

Answer by ewatrrr(24785)


Hi
normally distributed:  µ = 65 and σ = 9
Using a TI 0r similarly an inexpensive calculator like Casio fx-115 ES plus:
Continuous curve:
P(x < 54) = P(x ≤ 54) =  normalcdf(-9999,54,65,9) = .1108
P(x> 80) = 1 - P(x ≤ 80) = 1 - normalcdf(-9999,80,65,9) = 1- .9522 = .1478
P(70 < x < 86) = normalcdf(70,86,65,9) = .2794

Important You feel comfortable using Your calculator
and knowing it can be used for in-class testing.
Wish You the Best in your Studies.